the sum of three consecutive odd integers is 111
if n represents the least of the three integers , which of the following equations represents the statement above ?

A 3n = 111
B 3n+2 = 111
C 3n+ 4 = 111
D 3n+6 = 111
E 3n+9 = 111

Found 2 solutions by CharlesG2, Alan3354:
Answer by CharlesG2(834)

"the sum of three consecutive odd integers is 111
if n represents the least of the three integers , which of the following equations represents the statement above ? "
A 3n = 111
n = 111/3 = 37
37 + 39 + 41 = 90 + 7 + 9 + 11
90 + 7 + 9 + 11 = 90 + 16 + 11
90 + 16 + 11 = 106 + 11
106 + 11 = 117
117 does not equal 111, FALSE
B 3n+2 = 111
3n = 109
n = 36 1/3
n is not an integer, FALSE
C 3n+ 4 = 111
3n = 107
n = 35 2/3
n is not an integer, FALSE
D 3n+6 = 111
3n = 105
n = 35
35 + 37 + 39 = 90 + 5 + 7 + 9
90 + 5 + 7 + 9 = 90 + 12 + 9
90 + 12 + 9 = 102 + 9
102 + 9 = 111
111 does equal 111, TRUE
E 3n+9 = 111
3n = 102
n = 34
n is not odd, FALSE
Answer is D



Answer by Alan3354(69443)

n + n+2 + n+4 = 111
3n + 6 = 111